### v1

Abstract
Whole-slide images (WSIs) are critical for cancer diagnosis due to their ultra-high resolution and rich semantic content. However, their massive size and the limited availability of fine-grained annotations pose substantial challenges for conventional supervised learning. We propose DSAGL (Dual-Stream Attention-Guided Learning), a novel weakly supervised classification framework that combines a teacher-student architecture with a dual-stream design. DSAGL explicitly addresses instance-level ambiguity and bag-level semantic consistency by generating multi-scale attention-based pseudo labels and guiding instance-level learning. A shared lightweight encoder (VSSMamba) enables efficient long-range dependency modeling, while a fusion-attentive module (FASA) enhances focus on sparse but diagnostically relevant regions. We further introduce a hybrid loss to enforce mutual consistency between the two streams. Experiments on CIFAR-10, NCT-CRC, and TCGA-Lung datasets demonstrate that DSAGL consistently outperforms state-of-the-art MIL baselines, achieving superior discriminative performance and robustness under weak supervision.
keywords:
[…]
{highlights}
We propose DSAGL, a novel weakly supervised classification framework that integrates a dual-stream structure and a teacher–student mechanism to jointly enhance instance-level and bag-level performance.
An alternating training strategy is introduced to improve semantic consistency and enable effective collaboration between the teacher and student branches.
We design a lightweight encoder (VSSMamba) and a scale-aware attention module (FASA) to balance efficient long-range modeling and focus on diagnostically critical regions.
DSAGL consistently outperforms representative MIL-based methods on both synthetic and real-world pathological datasets at the instance and bag levels.
1 Introduction
In the field of digital medicine, ultra-high-resolution whole slide images (WSIs) have become a core imaging modality in modern pathological diagnosis due to their comprehensive and information-rich representation [35, 42, 8, 50]. WSIs can capture microscopic tissue structures with exceptional precision, providing critical support for cancer diagnosis and treatment planning. However, their massive size imposes substantial computational burdens [11, 43]. The complexity of tissue structures further complicates the accurate identification of abnormal regions, which often exhibit diffuse and scattered morphological features that hinder automated analysis. Traditional supervised learning methods rely heavily on exhaustive annotations by expert pathologists—a process that is not only time-consuming and costly but also subject to significant inter-observer variability, ultimately compromising model generalization. To address these challenges, weakly supervised learning approaches have gained increasing attention in computational pathology [27, 29, 34, 38, 44]. By leveraging easily accessible patient-level labels, these methods reduce the dependency on high-quality annotations and offer a promising pathway toward efficient and accurate analysis of large-scale pathological images.
In recent years, significant progress has been made in whole-slide image (WSI) classification for computational pathology, achieving notable progress. Among these approaches, multiple instance learning (MIL) [18, 19, 26, 31] has emerged as a dominant paradigm. MIL-based methods typically regard a WSI as a “bag” composed of numerous image patches and leverage patient-level labels to guide instance-level learning [51, 3]. This formulation not only makes effective use of coarse-grained annotations but also aligns well with clinical diagnostic practices, significantly reducing the cost of data labeling while delivering promising results across various tasks. However, existing MIL methods often rely on the assumption that instances within a bag are independent and identically distributed (i.i.d.), or adopt simplistic aggregation strategies such as attention-based pooling [19]. These approaches are limited in their ability to capture complex spatial dependencies and contextual information that are critical for accurately identifying diagnostic regions. The lack of long-range modeling across spatially distant patches hinders precise localization and recognition of relevant lesions, especially in real-world scenarios characterized by sparse and heterogeneous pathological patterns. Moreover, a fundamental trade-off persists between instance-level discriminability and bag-level classification performance, which compromises the model’s generalization and clinical applicability. Therefore, under weak supervision, designing more expressive and context-aware architectures that can jointly enhance instance-level and bag-level performance remains a central challenge in computational pathology.
Despite significant progress in improving diagnostic accuracy and computational efficiency for WSI classification, current methods still face two fundamental challenges that severely limit their robustness and generalization capabilities. First, most existing approaches rely on convolutional neural networks (CNNs) [1] or Transformer-based architectures [14, 30]. CNNs are constrained by their limited receptive fields, making it difficult to effectively model long-range dependencies [12, 17, 53]. Transformers, while capable of capturing global context, suffer from quadratic computational complexity with respect to the number of instances, resulting in poor scalability and inefficiency in high-resolution WSI scenarios [53, 17]. Second, conventional attention mechanisms tend to focus on densely textured or visually consistent regions, often overlooking sparse yet diagnostically critical instances. This issue is particularly pronounced in widely used single-stream MIL architectures [20], which lack explicit mechanisms for semantic refinement and instance-level disambiguation. As a result, such models often struggle to accurately discriminate key features at the instance level, leading to reduced robustness and generalization in the presence of complex and heterogeneous lesions (see Fig. 1(a)). To address these limitations, the emergence of Mamba offers a promising new approach for modeling long-range dependencies. By enabling efficient global context modeling with linear computational complexity, Mamba significantly reduces the processing cost of large-scale instance sets without compromising performance.
To address the aforementioned challenges, we propose DSAGL (Dual-Stream Attention-Guided Learning), a novel weakly supervised classification framework based on a dual-stream, teacher–student architecture with attention guidance. DSAGL is designed to capture inter-instance correlations while jointly enhancing both instance-level discriminability and bag-level classification performance in MIL models. The framework consists of a unified feature extraction backbone and a dual-stream structure. Specifically, the teacher stream employs a multi-scale attention mechanism to focus on diagnostically relevant instances, enabling accurate bag-level classification while generating high-quality soft pseudo labels. Guided by these pseudo labels, the student stream performs instance-level classification, thereby improving fine-grained discriminative capability. Moreover, as shown in Fig. 1(b), DSAGL employs an alternating training strategy between the teacher and student networks, in contrast to the sequential paradigm adopted by conventional dual-stream models. This collaborative optimization enhances semantic consistency and further improves the overall model performance.
Figure 1: (a) Basic MIL architecture; (b) Our proposed DSAGL architecture.
To further enhance the model’s representational capacity and generalization under weak supervision, DSAGL incorporates two key components: a Vision-aware Skip-connected Selective Mamba encoder (VSSMamba) for efficient long-range dependency modeling with linear complexity, and a Fusion-Attentive Scale-Aware module (FASA) to emphasize sparse but diagnostically relevant regions. These designs enable DSAGL to effectively leverage coarse-grained labels for accurate WSI classification. Extensive experiments conducted on both two synthetic datasets (CIFAR-10 and NCT-CRC) and one real-world pathological dataset (TCGA-Lung) demonstrate that DSAGL outperforms representative MIL-based approaches at both instance and bag levels, highlighting its strong discriminative power and generalization across diverse scenarios.
The main contributions of this work are summarized as follows:
  • We propose a novel weakly supervised classification framework, DSAGL, which integrates a dual-stream architecture with a teacher–student mechanism to jointly optimize instance-level discriminability and bag-level classification performance, addressing the limitations of conventional MIL methods in semantic modeling.
  • We introduce an alternating training strategy to enhance semantic consistency, enabling collaborative optimization between the teacher and student branches, which significantly improves the model’s robustness and generalization in complex lesion recognition scenarios.
  • A lightweight backbone VSSMamba for efficient long-range dependency modeling and a multi-scale attention module FASA for capturing discriminative patterns in sparse tumor regions.
The remainder of this paper is organized as follows: Section 2 reviews related work and highlights current limitations and challenges. Section 3 focuses on the proposed model architecture, detailing its design and key technical components. Section 4 presents experimental results on three datasets along with ablation studies. Section 5 discusses the strengths and limitations of the proposed approach. Finally, Section 6 concludes the paper and outlines potential directions for future research.
2 Related Work
2.1 Multiple Instance Learning for WSI Classification
Multiple instance learning (MIL) [47, 40, 40] has become a widely adopted weakly supervised paradigm in computational pathology, particularly well-suited for whole slide images (WSIs), which are characterized by extremely high resolution and large-scale tissue content. In MIL, each WSI is divided into a set of patches (instances) grouped into a bag, with supervision provided only at the bag level. This eliminates the need for dense annotations and significantly reduces manual labeling costs.
Early MIL methods, such as MI-Net [48], employed max or mean pooling to aggregate instance features for bag-level prediction. While computationally efficient, these approaches are limited in their ability to distinguish semantically relevant instances. To address this, Ilse et al. [19] introduced Attention-MIL, which incorporates learnable attention weights to assign importance scores to instances. The framework was further extended to ABMIL by incorporating gated attention mechanisms, thereby improving both aggregation flexibility and interpretability.  [19] Subsequent works have explored more expressive instance-level modeling strategies. Chen et al. [5] proposed MSA-MIL by integrating multi-scale feature extraction modules, enhancing sensitivity to fine-grained structural variations. Li et al. [28] introduced SA-MIL, which embeds self-attention into the aggregation process to capture contextual dependencies among instances. DTFD-MIL [51] employs a dual-path differential feature design to strengthen the model’s ability to detect heterogeneous lesion patterns. In addition, DSMIL [26] introduces an explicit instance selection module within a dual-branch structure, enabling collaborative instance- and bag-level decision-making. CLAM [31] incorporates attention pruning to localize diagnostically relevant regions and improve interpretability. Transformer-based frameworks such as TransMIL [39] further extend MIL by modeling global dependencies across instances, exhibiting superior long-range reasoning compared to traditional attention-based models.
Despite these advances, two key limitations remain in current MIL-based approaches. First, most models rely solely on bag-level supervision and lack explicit instance-level guidance, which hinders precise localization of critical pathological regions. Second, attention mechanisms are typically used only for feature aggregation and are not explicitly aligned with classification objectives, limiting their effectiveness under weak supervision.
2.2 Dual-Stream Architectures and Knowledge Distillation
Prior to the emergence of dual-stream architectures, extensive research explored the integration of knowledge distillation (KD) into single-branch or loosely coupled frameworks to enhance the representational capacity and training stability of student models. For example, Co-MIL [20] employs a multi-teacher collaborative distillation strategy, where separate teacher models extract global and local semantic information to guide the student through cross-model supervision. GLMC [10] introduces bidirectional guidance between global and local branches, thereby improving the model’s ability to capture multi-scale lesion features. Self-Distillation utilizes the historical states of the student model as an implicit teacher, enabling knowledge transfer without additional parameter overhead—an approach particularly suitable for resource-constrained settings. ReNA-KD [52] incorporates region-structured attention maps into the distillation process, enhancing spatial awareness and improving recognition of heterogeneous lesion patterns.
Building upon these foundational works, recent studies have further integrated knowledge distillation into dual-stream frameworks to enhance semantic consistency and cross-branch collaboration. R2T-MIL [45] introduces a region-aware token distillation module within a Transformer-based architecture, leveraging spatial attention to guide the student in focusing on diagnostically relevant areas. GLMKD [6] designs a dual-teacher framework with global and local branches, and incorporates a shape transfer loss to reconcile inconsistencies between pseudo labels from different teachers, enabling more accurate boundary modeling and improved training stability under weak supervision.
While recent methods have introduced improvements, many distillation strategies remain constrained by static soft labels and rigid optimization phases. The absence of adaptive feedback mechanisms hampers their capacity to cope with semantic drift and evolving inter-branch representations during training.
2.3 Long-Range Dependency Modeling in WSI
Whole Slide Images (WSIs) are composed of thousands of image patches and exhibit strong non-local structural characteristics. In pathological diagnosis, spatially distant regions often present semantic dependencies—for example, between tumor boundaries and cores, or between inflamed tissues and surrounding microenvironments. Modeling such long-range dependencies is critical for improving classification accuracy.
Convolutional neural networks (CNNs) [2, 7] are inherently limited by their local receptive fields and struggle to capture global contextual information. To address this, Transformer-based [46, 15] architectures have been introduced. TransMIL [39] is one of the earliest attempts to model inter-patch relationships via self-attention for global bag representation. Vision Transformer (ViT) [9] extends standard Transformer designs to image modeling and has been widely adopted in MIL settings. DTFD-MIL further incorporates feature differencing to enhance the modeling of local-global disparities.
Despite their expressive capacity, Transformer models suffer from high computational cost. The attention mechanism scales quadratically with input size, making it inefficient for WSIs with thousands of patches. This leads to memory bottlenecks and latency issues, hindering real-world deployment. To address these limitations, Structured State Space Models (SSMs) [13] have been proposed as efficient alternatives. S4, introduced by Gu et al., first incorporated continuous-time state-space equations into deep networks for efficient long-range sequence modeling. Building on this, Mamba introduced selective state updates and input gating mechanisms, enabling fine-grained control over information flow while preserving the efficiency of SSMs. Mamba [12, 17, 32] achieves superior parallelism and scalability, outperforming standard Transformers in tasks such as image classification and semantic segmentation.
Derived from the Mamba framework, several variants have been developed. Vision Mamba [53] integrates convolutional pre-processing and spatial shift mechanisms to enhance spatial awareness. T-Mamba [17] introduces temporal decoupling for efficient modeling of non-uniform sequential data. U-Mamba [32] extends the architecture for biomedical image segmentation by incorporating a U-shaped encoder–decoder design with Mamba blocks, enabling long-range context modeling while preserving spatial resolution.
3 Method
This section provides a detailed exposition of the proposed methodological framework, encompassing the formal definition of the problem, the overall architectural design of the model, and the construction principles of key components, including the VSSMamba and the Dual-Stream Network.
3.1 Task Definition
3.1.1 Multiple Instance Learning (MIL)
We consider weakly supervised pathological image classification under the Multiple Instance Learning (MIL) framework. As illustrated in Fig. 2, unlike traditional supervised learning where each instance is individually labeled, MIL assigns labels only at the bag level, with each bag containing multiple unlabeled instances.
Each Whole Slide Image (WSI) is divided into a bag of image patches, where each patch is treated as an instance [19, 27, 29, 41]. A bag is defined as $B_{i}=\{x_{i,1},\dots,x_{i,N_{i}}\}$, where $x_{i,j}\in\mathbb{R}^{\texttt{d}}$ denotes the d-dimensional feature vector of the $j$-th instance in bag $i$, and $N_{i}$ is the number of instances.
The dataset is represented as $\mathcal{D}=\{(B_{i},y_{i})\}_{i=1}^{M}$, where only bag-level labels $y_{i}\in\{0,1\}$ are available during training. According to the standard MIL assumption, a bag is positive if it contains at least one positive instance:
$y_{i}=\max_{1\leq j\leq N_{i}}z_{i,j},$ (1)
where $z_{i,j}\in\{0,1\}$ denotes the (latent) instance label. The goal is to jointly optimize bag-level classification and instance-level discrimination under weak supervision.
Figure 2: Supervised learning and Multiple-Instance Learning (MIL).
3.1.2 SSM and Mamba
Structured State Space Models (SSMs) [53] offer an efficient module for long-range sequence modeling with linear time complexity. Given a 1D input sequence $x(t)$, the system evolves through a hidden state $h(t)\in\mathbb{R}^{\texttt{N}}$, where N denotes the dimension of the hidden state, governed by the continuous-time dynamics:
$\frac{d}{dt}h(t)=\mathbf{A}h(t)+\mathbf{B}x(t),\quad y(t)=\mathbf{C}h(t),$ (2)
where $\mathbf{A}\in\mathbb{R}^{\texttt{N}\times\texttt{N}},\mathbf{B}\in\mathbb{R}^{\texttt{N}\times\texttt{1}},\mathbf{C}\in\mathbb{R}^{\texttt{1}\times\texttt{N}}$ are learnable matrices. For discrete-time processing, we apply zero-order hold (ZOH) with time step $\Delta$, yielding:
$\mathbf{\overline{A}}=\exp(\Delta\mathbf{A})$ (3)
$\mathbf{\overline{B}}=\left(\Delta\mathbf{A}\right)^{-1}\left(\exp(\Delta\mathbf{A})-\mathbf{I}\right)\Delta\mathbf{B}$ (4)
$h_{t}=\mathbf{\overline{A}}h_{t-1}+\mathbf{\overline{B}}x_{t},\quad y_{t}=\mathbf{C}h_{t}$ (5)
The output sequence can also be expressed as a global convolution over a structured kernel:
$\mathbf{\overline{K}}=\left(\mathbf{C}\mathbf{\overline{A}}^{i}\mathbf{\overline{B}}\right)_{i=0}^{L-1}$ (6)
$y=\mathbf{\overline{K}}\otimes x$ (7)
where $\otimes$ denotes one-dimensional convolution. Mamba [12] implements this computation efficiently using either convolution or recurrence, and further enhances it with input-dependent parameterization, enabling scalable modeling for long sequences in applications such as language and video understanding.
3.2 Overall Framework
3.2.1 DSAGL architecture
Figure 3: (a) The overall architecture of the proposed Dual-Stream Attention-Guided Learning Weakly Supervised Framework (DSAGL); (b) The proposed Vision-aware Skip-connected Selective Mamba encoder (VSSMamba); (c) The proposed Fusion-Attentive Scale-Aware module (FASA).
To achieve efficient and interpretable pathological image classification, we propose a Dual-Stream Attention-Guided Learning Weakly Supervised Framework (DSAGL). As illustrated in Fig. 3, the DSAGL architecture consists of two main components: a feature extraction module and a dual-stream network comprising a teacher branch and a student branch. Specifically, we first employ the Vision-aware Skip-connected Selective Mamba encoder (VSSMamba) to extract preliminary instance-level features from the input images. These features are then processed independently by the teacher and student branches. In the teacher branch, we introduce a Fusion-Attentive Scale-Aware module (FASA) to generate bag-level predictions and compute attention scores, which are subsequently used to construct high-quality instance-level pseudo-labels. The student branch focuses on instance-level classification, using the pseudo-labels from the teacher as supervision signals. By alternately updating the teacher and student networks, DSAGL facilitates efficient and collaborative feature sharing and representation learning.
3.2.2 Alternating Training Strategy
To enhance the collaborative optimization between the teacher and student networks under weak supervision, DSAGL adopts an alternating training strategy. Rather than employing conventional joint or sequential training paradigms, this strategy updates the two branches in an interleaved manner, enabling iterative refinement of pseudo labels and feature representations throughout the training process. During training, the teacher network is updated at fixed intervals, while the student network is optimized continuously using the pseudo labels generated by the teacher. This staged update pattern mitigates potential gradient interference between the two branches and contributes to stable convergence. Moreover, by decoupling the training dynamics, each network can specialize in its respective task: global bag-level prediction for the teacher and fine-grained instance-level discrimination for the student.
This alternating strategy offers several key advantages. First, it facilitates dynamic refinement of pseudo labels, which are progressively improved throughout training, thereby reducing the impact of noise in the early stages. Second, it enables a clear decoupling of tasks between the two branches, allowing the teacher to focus on holistic aggregation while the student learns detailed discriminative features. Third, the staggered updates improve overall training stability by avoiding simultaneous gradient conflicts. Finally, this design proves particularly effective in weakly supervised scenarios, where sparse or noisy labels can severely degrade performance.
3.3 VSSMamba (Feature Extraction)
Although CNNs [25] and Transformers [46] have excelled in vision tasks, they remain limited in modeling long-range dependencies and incur high computational costs on large-scale WSIs.To address these limitations, inspired by Vision Mamba [53], we propose VSSMamba, a new encoder that extracts discriminative instance-level features with linear complexity.
As illustrated in part (b) of Fig. 3, the VSSMamba comprises a main processing stream and two auxiliary skip connections, enabling enhanced multi-scale and multi-level context fusion. Given an input instance image $p\in\mathbb{R}^{\texttt{B}\times\texttt{C}\times\texttt{H}\times\texttt{W}}$, where B, C, H, and W denote the batch size, number of channels, height, and width respectively, we first extract initial features via a convolutional stem layer:
$f_{0}=\mathrm{Stem}(p)$ (8)
The features are subsequently processed by a stack of $n$ Mamba layers, denoted as $\{\mathcal{M}_{i}\}_{i=1}^{n}$, along the main path. Each layer consists of a Mamba transformation, followed by Dropout, Efficient Channel Attention $\mathcal{E}_{i}$, and a downsampling operation $\mathcal{D}_{i}$. The number of stacked layers $n$ can be flexibly adjusted based on data scale and task complexity to balance representation depth and capacity. We denote the final output of the main path as $f_{n}$ and output $f_{i+1}$ of the $i$-th layer is computed as:
$f_{i+1}=\mathcal{D}_{i}\left(\mathcal{E}_{i}\left(\mathrm{Drop}_{i}\left(\mathcal{M}_{i}(f_{i})\right)\right)\right),\quad i=0,1,\dots,n{-}1$ (9)
To enhance both local detail and global semantics, VSSMamba incorporates two skip connections, $f_{skip1}$ and $f_{skip2}$. The first skip path starts from the stem output $f_{0}$, and passes through a combined attention module consisting of Squeeze-and-Excitation and Spatial Attention $\mathcal{A}$. The output is then downsampled and regularized via Dropout to yield locally enhanced features. The second skip path directly downscales the raw input $p$ to extract global semantic cues:
$f_{skip1}=\mathrm{Drop}\left(\mathcal{D}\left(\mathcal{A}(f_{0})\right)\right)$ (10)
$f_{skip2}=\mathrm{Drop}\left(\mathcal{D}(p)\right)$ (11)
Finally, VSSMamba integrates the outputs from all three paths—$f_{n}$, $f_{skip1}$, and $f_{skip2}$ through a residual aggregation strategy, effectively blending deep semantic features, localized attention-enhanced cues, and global contextual signals. This fusion yields the final instance-level representation $x$, which captures both fine-grained structural detail and long-range contextual dependencies crucial for accurate instance discrimination in WSI:
$x=f_{n}+f_{skip1}+f_{skip2}$ (12)
In summary, VSSMamba efficiently captures both global dependencies and local details through its hierarchical design with linear-time complexity. By combining the long-range modeling ability of Mamba layers with complementary skip-enhanced feature streams, it provides robust and expressive instance-level representations. This encoder serves as the backbone of our dual-stream framework, enabling effective learning under weak supervision for complex WSI classification tasks.
{algorithm}
[H] VSSMamba Encoding Process Input: Instance patch $p:{\color[rgb]{0,0.46875,0}\mathtt{(B,C,H,W)}}$, number of layers $n$
[…]
$x$
3.4 Duel-Stream Network
To jointly model instance-level and bag-level semantics under weak supervision, we propose a dual-stream architecture comprising a teacher and a student network. This design is motivated by the inherent ambiguity in instance labels within WSIs, where relying solely on bag-level supervision often leads to suboptimal instance discrimination. By decoupling the representation learning into two cooperative branches, the teacher can provide global guidance, while the student focuses on fine-grained instance-level feature extraction.
In this section, we first present FASA, the key attention module designed to enhance the model’s ability to extract discriminative features from sparse and ambiguous regions. FASA integrates multi-scale convolution and dual attention mechanisms to address the challenge of lesion heterogeneity and low signal-to-noise ratio. We then elaborate on the functional roles and collaborative interaction between the teacher and student branches, followed by a detailed formulation of the hybrid loss function. This loss is tailored to mitigate the noise and uncertainty introduced by pseudo labels, thus promoting stable optimization and semantic alignment between the two streams.
3.4.1 FASA
To address the challenge of limited supervision, we introduce FASA, a fusion-attentive scale-aware module that enhances feature discriminability by integrating multi-scale convolutional fusion with both channel and spatial attention mechanisms. It performs attention-weighted global pooling to aggregate instance-level features, thereby facilitating discriminative learning in a weakly supervised setting.
As illustrated in part (c) of Fig. 3, FASA improves the discriminability of deep features by integrating spatial and channel attention into a multi-scale fusion framework. Given an input feature map $x\in\mathbb{R}^{\texttt{B}\times\texttt{C}\times\texttt{H}\times\texttt{W}}$ extracted by the VSSMamba encoder, FASA captures local contextual information at multiple receptive fields through parallel convolutions with kernel sizes of $3\times 3$, $5\times 5$, and $7\times 7$, where $\mathcal{C}_{k\times k}$ denotes a convolution operation of size $k\times k$. The resulting multi-scale features are subsequently concatenated and aggregated along the channel dimension, yielding a unified representation $h_{fused}$ that preserves complementary information across different spatial scales.
Next, channel and spatial attention mechanisms are independently applied to $h_{fused}$, producing two complementary attention-enhanced feature maps: $f_{channel}$ and $f_{spatial}$. These are then combined to obtain the final attentive representation $f_{out}$, which integrates both modality-specific importance cues. During instance aggregation, $f_{out}$ is subjected to global pooling to derive instance-wise attention weights $a_{i,j}$ and the corresponding bag-level prediction score $\text{logit}_{i}$. These outputs jointly serve as weak supervision signals for pseudo-label generation and bag-level classification.
The flow of feature computation is formulated as follows:
$h_{fused}=\sum_{k\in\{3,5,7\}}\mathcal{C}_{k\times k}(x),\quad k\in\{3,5,7\}$ (13)
$f_{t}=h_{fused}\times\mathcal{A}_{t}(h_{fused}),\quad t\in\{channel,spatial\}$ (14)
$f_{out}=f_{channel}+f_{spatial}$ (15)
$a_{i,j},\,logit_{i}=\text{FASA}(f_{out})$ (16)
By unifying multi-scale feature extraction with spatial and channel-wise attention, FASA provides a powerful mechanism for emphasizing diagnostically relevant regions under weak supervision. Its design allows the network to adaptively aggregate informative patterns across varying spatial resolutions, thereby reinforcing instance-level discriminability and improving the quality of attention-driven pseudo labels essential for downstream tasks.
keywordcolorRGB0, 120, 0 {algorithm}[H] FASA Block Process Input: Feature map $x:{\color[rgb]{0,0,0}\mathtt{(B,C,H,W)}}$
[…]
$a_{i,j},\textit{logit}_{i}$
3.4.2 Teacher Network
The teacher network serves as a bag-level classifier that provides both global predictions and instance-wise guidance under weak supervision. This design enables the decoupling of coarse-grained supervision from fine-grained feature learning, allowing the teacher to learn stable bag-level semantics while generating informative cues for the student network.
Upon receiving instance-level feature vectors, the teacher network employs FASA to compute attention scores for each instance and predicts the bag label via a classification head. The attention mechanism allows the teacher to focus on diagnostically relevant patches and suppress irrelevant or noisy regions, which is particularly important in WSIs where discriminative signals are sparse and spatially scattered.
Specifically, for each instance $\left\{p_{i,j},\,j=1,2,\dots,n_{i}\right\}$ in a bag $B_{i}$, the corresponding feature vector $x_{i,j}$ is passed into FASA to obtain its attention score $a_{i,j}$ and the aggregated bag-level logit $logit_{i}$. The attention scores $\{a_{i,j}\}$ are propagated to the student network as soft supervisory signals, while the logit is further processed by the classification head $P_{t}$ to generate the predicted bag label $\hat{y}_{i}$. The bag prediction is supervised using the standard binary cross-entropy loss $\mathcal{L}_{teacher}$, which is robust and widely adopted for binary weakly labeled classification tasks. The computation proceeds as follows:
$a_{i,j},\,logit_{i}=\text{FASA}\left(x_{i,j}\mid x_{i,1},x_{i,2},\dots,x_{i,n_{i}}\right)$ (17)
$\hat{y}_{i}=P_{t}(logit_{i})$ (18)
$\mathcal{L}_{T}=\mathcal{L}_{teacher}(y_{i},\hat{y}_{i})$ (19)
By decoupling semantic abstraction and region selection, the teacher network delivers reliable bag-level predictions while producing spatially-aware attention signals. These outputs play a pivotal role in guiding the student network, particularly under noisy and weakly annotated scenarios common in WSI tasks.
3.4.3 Student Network
The student network is designed to perform fine-grained instance-level classification under weak supervision. It receives instance feature vectors and predicts the class probability for each individual instance. While sharing the same feature extractor as the teacher network, the student maintains its own classification head $P_{s}$, enabling independent optimization of instance-level representations.
To guide the student in identifying diagnostically relevant regions, the attention score $a_{i,j}$ provided by the teacher is passed through a normalization module, NormProb, which converts it into a probabilistic pseudo label $\hat{z}_{i,j}$. This process stabilizes the supervision signal and helps mitigate early-stage noise in the attention distribution. The training of the student network is supervised using a hybrid loss function $\mathcal{L}_{Hybrid}$, which combines cross-entropy loss [33] with Kullback–Leibler divergence [24] in a weighted formulation. This design allows the network to benefit from both hard pseudo labels and soft distribution alignment. The resulting student loss is denoted by $\mathcal{L}_{S}$, and is computed as follows:
$\hat{z}_{i,j}=\text{NormProb}(a_{i,j})$ (20)
$p_{\hat{z}_{i,j}}=P_{s}(x_{i,j})$ (21)
$\mathcal{L}_{S}=\mathcal{L}_{Hybrid}(\hat{z}_{i,j},p_{\hat{z}_{i,j}})$ (22)
By leveraging attention-informed pseudo labels and a tailored hybrid loss, the student network effectively refines instance-level predictions under weak supervision. Its design enables focused learning on diagnostically salient regions and complements the teacher branch by reinforcing semantic consistency at a finer granularity.
3.4.4 Hybrid Loss Function
Under weak supervision, instance-level labels are unavailable, and pseudo labels generated by the teacher network are inherently noisy. Relying solely on cross-entropy (CE) loss can lead to overfitting and unstable optimization, especially in early training stages. To address this, we propose a hybrid loss function $\mathcal{L}_{Hybrid}$ that combines weighted hard-label supervision with soft-label regularization, balancing discriminative learning and robustness.
The student network is supervised by two complementary objectives: a weighted cross-entropy loss $\mathcal{L}_{CE,student}$ that promotes confident classification based on hard pseudo labels, and a Kullback-Leibler (KL) divergence loss $\mathcal{L}_{KL,student}$ that aligns the student’s output distribution with the teacher’s soft predictions. This dual supervision enables sharper decision boundaries while preserving probabilistic consistency. To mitigate instance-level class imbalance, a weight factor $w_{n}$ is applied to negative samples in the CE loss.
Formally, the teacher network is optimized with a bag-level cross-entropy loss:
$\mathcal{L}_{teacher}=-\left[y_{i}\log(\hat{y}_{i}+\epsilon)+(1-y_{i})\log(1-\hat{y}_{i}+\epsilon)\right],$ (23)
where $y_{i}$ denotes the ground-truth bag label, $\hat{y}_{i}$ is the predicted bag-level probability, and $\epsilon$ is a small constant for numerical stability.
The student cross-entropy loss is defined as:
$\mathcal{L}_{CE,student}=-\frac{1}{N}\sum_{i,j}\left[w_{n}(1-\hat{z}_{i,j})\log(p_{i,j,0}+\epsilon)+(1-w_{n})\hat{z}_{i,j}\log(p_{i,j,1}+\epsilon)\right],$ (24)
where $\hat{z}_{i,j}\in\{0,1\}$ is the hard pseudo label for instance $(i,j)$, $p_{i,j,c}$ is the student’s predicted probability for class $c\in\{0,1\}$, and $w_{n}\in[0,1]$ controls the relative importance of negative instances.
To leverage soft supervision, the KL divergence loss is introduced as:
$\mathcal{L}_{KL,student}=T^{2}\cdot\frac{1}{N}\sum_{i,j}\sum_{c\in\{0,1\}}q_{i,j,c}\log\left(\frac{q_{i,j,c}}{p_{i,j,c}}\right),$ (25)
where $q_{i,j,c}$ and $p_{i,j,c}$ are the predicted class probabilities from the student and teacher, respectively. The temperature parameter $T>1$ is used to soften the distributions, and the KL loss is scaled by $T^{2}$ to match the magnitude of CE gradients.
The overall student loss is expressed as:
$\mathcal{L}_{Hybrid}=\alpha\mathcal{L}_{CE,student}+(1-\alpha)\mathcal{L}_{KL,student},$ (26)
where $\alpha\in[0,1]$ controls the trade-off between hard and soft supervision. The hybrid loss effectively stabilizes learning and improves generalization, particularly under noisy pseudo labels and imbalanced patch distributions.
4 Experiment and Results
4.1 Datasets
To comprehensively evaluate the performance of DSAGL, we employed three datasets of increasing image resolution, encompassing both synthetic and real-world scenarios, as illustrated in Fig. 4.
Figure 4: Examples of the three datasets used in our experiments. (a) CIFAR-10, consisting of 10 categories of natural images; (b) NCT-CRC, consisting of 9 categories of colorectal histopathology image patches; (c) TCGA-Lung, consisting of whole-slide images from two lung cancer subtypes.
4.1.1 Cifar-10
Cifar-10 [23] is a standard image classification dataset consisting of 60,000 color images of size $32\times 32$ pixels, divided into 10 classes, with 6,000 images per class. Among them, 50,000 images are used for training and 10,000 for testing. The training set is organized into five batches of 10,000 images each, while the test set constitutes a separate batch. In weakly supervised learning tasks, Cifar-10 is commonly utilized to simulate whole-slide images (WSIs) by randomly assembling images from different classes to construct synthetic datasets. In our setting, we simulate WSIs by randomly combining images from different classes within the dataset.
4.1.2 NCT-CRC
NCT-CRC-HE-100K [21] is a real-world histopathology image dataset comprising 100,000 non-overlapping image patches extracted from 86 H&E-stained human colorectal cancer and normal tissue slides. All images are of size $224\times 224$ pixels and sampled at 20×20$\times$20 × magnification. Each image is annotated with one of nine histological tissue classes: adipose (ADI), background (BACK), debris (DEB), lymphocytes (LYM), mucus (MUC), smooth muscle (MUS), normal colon mucosa (NORM), cancer-associated stroma (STR), and colorectal adenocarcinoma epithelium (TUM). Based on this dataset, we constructed multiple subsets with varying positive instance ratios by selecting images from different tissue classes.
4.1.3 TCGA-Lung
The TCGA (The Cancer Genome Atlas) lung cancer dataset [26] is an important publicly available resource for histopathological image analysis, encompassing two major subtypes of non-small cell lung cancer: lung adenocarcinoma (LUAD) and lung squamous cell carcinoma (LUSC). The dataset comprises a total of 1,054 high-resolution whole-slide images (WSIs) stained with hematoxylin and eosin (H&E), provided in the svs format and accessible through the data portal of the National Cancer Institute of the United States. At 20×20$\times$20 × magnification, each WSI is partitioned into approximately 5.2 million image patches of size $224\times 224$ pixels. Each WSI constitutes a bag in the multiple instance learning setting, containing a varying number of tumor and normal patches.
4.2 Experimental Details
In the experimental setup, we observe that the instance resolution in the Cifar-10 dataset is relatively low, which makes it less suitable for the Mamba architecture designed for modeling long-range dependencies. Consequently, the number of Mamba layers along the main path (as illustrated in Fig. 3) is set to zero, while the convolutional depth of the Stem module is increased to enhance feature extraction capability. In contrast, for the NCT-CRC and TCGA-Lung datasets, where the instance resolution is higher, the number of Mamba layers is set to 2 and 3, respectively.
In order to facilitate more stable gradient accumulation and efficient parameter updates, we employ the Adam optimizer [22] on NCT-CRC, with a fixed learning rate of 0.0001 and a weight decay of $e^{-5}$, and set the batch size to 4. For Cifar-10 and TCGA-Lung, we adopt the SGD optimizer [37] with a learning rate of 0.001 and a batch size of 1. All experiments are conducted on a single NVIDIA V100 GPU.
For both instance-level and bag-level classification tasks, we adopt the area under the ROC curve (AUC) [16] as the primary evaluation metric. Specifically, we report the AUC scores of the DSAGL on both instance-level and bag-level tasks for the synthetic datasets. However, since the TCGA-Lung cancer dataset does not provide ground-truth instance-level labels, we only report the bag-level AUC for this dataset.
Table 1: Results on the CIFAR-10 dataset.
(a) Instance-level classification AUC.
      Positive patch ratio       1%       5%       10%       20%       50%       70%
      MI-Net [48]       0.6447       0.7565       0.7803       0.7933       0.7822       0.7674
      Attn-MIL [19]       0.6225       0.6213       0.7370       0.7551       0.7028       0.6342
      Gated-Attn-MIL [19]       0.6187       0.6873       0.7970       0.7831       0.6378       0.6277
      ABMIL [19]       0.6253       0.9083       0.9241       0.9237       0.8224       0.7935
      DSMIL [26]       0.4039       0.5515       0.4918       0.8258       0.6152       0.7525
      DSAGL       0.5971       0.8552       0.9037       0.9438       0.9092       0.9103
(b) Bag-level classification AUC.
      Positive patch ratio       1%       5%       10%       20%       50%       70%
      MI-Net [48]       0.6114       0.6952       0.7388       0.8640       0.9075       1.000
      Attn-MIL [19]       0.6235       0.6589       0.8832       1.000       1.000       1.000
      Gated-Attn-MIL [19]       0.5444       0.5792       0.8637       1.000       1.000       1.000
      ABMIL [19]       0.5783       0.8850       0.9955       1.000       1.000       1.000
      DSMIL [26]       0.4025       0.5174       0.5265       0.9468       0.9850       1.000
      DSAGL       0.6044       0.8203       0.9715       1.000       1.000       1.000
4.3 Comparison Methods
We compare DSAGL with a diverse set of representative multiple instance learning (MIL) approaches, encompassing classical pooling-based and attention-based models (e.g., MI-Net [48], Attention-MIL [19], ABMIL [19]), as well as more recent architectures that incorporate dual-stream mechanisms, Transformer-based global modeling, and region-aware learning strategies, including DSMIL [26], CLAM [31], TransMIL [39], IBMIL [29], and R2T-MIL [45]. For a fair comparison, both ABMIL and DSMIL adopt SimCLR [26] as their feature extractor. Experiments are conducted on the Cifar-10 and NCT-CRC datasets under varying positive instance ratios ranging from 10% to 70%. As the TCGA-Lung dataset lacks ground-truth instance-level annotations, only bag-level evaluation is performed. For several MIL baselines, part of the results are reproduced in our experimental settings, while others are directly referenced from existing studies on weakly supervised WSI classification [36].
4.4 Experimental Results
For synthetic datasets, Table 1 and Table 2 report the instance-level and bag-level classification performance of DSAGL on CIFAR-10 and NCT-CRC under varying positive patch ratios. As shown in Table 1, DSAGL exhibits strong stability across different settings, achieving instance-level AUCs of 0.9438, 0.9092, and 0.9103 at 20%, 50%, and 70%, respectively—outperforming baselines such as ABMIL and DSMIL. At the bag level, DSAGL achieves a perfect AUC of 1.000 across all three ratios, further confirming its robustness under high signal-to-noise conditions. These results demonstrate DSAGL’s ability to remain reliable in synthetic weak supervision scenarios and its resilience to noisy pseudo labels.
On the NCT-CRC dataset (Table 2), DSAGL consistently outperforms all competing methods across different positive ratios. Its instance-level AUC remains above 0.95 in all settings, reaching a peak of 0.9916 at 70%, highlighting its robustness to distributional shifts and morphological diversity. Additionally, DSAGL achieves a perfect bag-level AUC of 1.000 under all conditions, surpassing or matching the performance of state-of-the-art MIL methods. This consistent superiority confirms DSAGL’s capability to generalize across complex tissue patterns while maintaining high discriminative accuracy at both instance and bag levels.
Table 2: Results on the NCT-CRC dataset.
(a) Instance-level classification AUC.
Pos. ratio 10% 20% 50% 70%
MI-Net [48] 0.6648 0.6950 0.6996 0.6899
Attn-MIL [19] 0.6621 0.6424 0.6409 0.6548
Gated-Attn-MIL [19] 0.6753 0.6333 0.6830 0.7228
ABMIL [19] 0.7410 0.8729 0.8800 0.7965
DSMIL [26] 0.3690 0.7008 0.4835 0.7399
DSAGL 0.9731 0.9538 0.9755 0.9916
(b) Bag-level classification AUC.
Pos. ratio 10% 20% 50% 70%
MI-Net [48] 0.7160 0.5185 0.7600 0.7212
Attn-MIL [19] 0.7407 0.7037 1.000 1.000
Gated-Attn-MIL [19] 0.9506 0.9506 1.000 1.000
ABMIL [19] 0.9754 1.000 0.9766 0.9894
DSMIL [26] 1.000 0.8914 0.9997 1.000
DSAGL 1.000 1.000 1.000 1.000
For the real-world TCGA-Lung dataset, Table 3 summarizes the bag-level classification results. Compared to a wide range of representative methods, DSAGL achieves the highest AUC of 0.9727, outperforming CLAM, TransMIL, and the recent R2T-MIL. This strong performance further validates the generalizability and clinical applicability of DSAGL, demonstrating its practical utility in real-world histopathological diagnosis where annotation is limited and data variability is substantial.
4.5 Ablation Study
To systematically evaluate the contribution of each core component within the DSAGL to classification performance at different granularities, we conducted an ablation study on the NCT-CRC dataset with a positive instance ratio of 10%. The results are summarized in Table 4.
Table 3: Results on the TCGA-Lung dataset.
Method Bag-level AUC
Max-pooling [49] 0.9014
Mean-pooling [49] 0.9369
RNN-MIL [4] 0.9107
ABMIL [19] 0.9488
DSMIL [26] 0.9533
CLAM [31] 0.9607
TransMIL [39] 0.9672
IBMIL [29] 0.9689
R2T-MIL [45] 0.9701
DSAGL 0.9727
Table 4: Ablation study on the CRC-MIL dataset.
       Duel-stream        FASA        $\mathcal{L}_{Hybrid}$        VSSMamba        Instance AUC        Bag AUC
       0.7410        0.9754
       ✓        0.8021        1.000
       ✓        ✓        0.9214        1.000
       ✓        ✓        ✓        0.9348        1.000
       ✓        ✓        ✓        ✓        0.9731        1.000
Starting from a baseline model without any additional modules, the instance-level AUC is 0.7410, and the bag-level AUC is 0.9754. Introducing the dual-stream architecture improves the instance-level AUC to 0.8021, while the bag-level AUC reaches 1.000, indicating that the teacher-student collaborative optimization mechanism effectively enhances bag-level prediction performance. Incorporating the FASA further boosts the instance-level AUC to 0.9214, demonstrating the effectiveness of multi-scale saliency modeling in identifying lesion regions.
Building on this, the addition of the hybrid loss function $\mathcal{L}_{Hybrid}$ leads to a further increase in instance-level AUC to 0.9348, highlighting the robustness advantages of joint supervision through hard and soft pseudo labels in weakly supervised learning. Finally, integrating the customized feature encoder VSSMamba into the overall architecture yields the best performance, achieving an instance-level AUC of 0.9731 while maintaining a bag-level AUC of 1.000, validating the effectiveness of this module in modeling long-range dependencies and fusing multi-scale features.
4.6 Visualization and Analysis
To further interpret the classification behavior of DSAGL, we visualize the attention responses from the student branch. Specifically, Grad-CAM is applied to the final convolutional layer of the student encoder, with gradients back-propagated from the classification output to generate class-discriminative attention heatmaps. These visualizations, derived exclusively from the student network, reflect its instance-level decision preferences and provide insights into the spatial reasoning patterns learned during training.
As shown in Fig. 5, we present attention maps across four representative tissue categories in the CRC-MIL dataset. For ADI (adipose tissue), the model successfully suppresses large homogeneous regions and focuses on interstitial lines and localized boundaries, indicating robustness to non-informative areas. In MUS (muscle), attention highlights fiber boundaries and inter-fiber gaps, capturing structural cues effectively. For LYM (lymphocytes), the model accurately detects sparse lymphocyte clusters, demonstrating strong granular attention to small-scale features. In the TUM (tumor) category, attention is highly concentrated on densely packed epithelial structures with nuclear atypia and glandular distortion, reflecting the model’s capacity for complex morphological reasoning and multi-focal response.
Figure 5: Visualization of attention heatmaps on representative categories from the CRC-MIL dataset. Each pair shows the original image (left) and the corresponding attention heatmap (right) generated by the student model.
We also observe occasional failure cases. In certain MUS samples, attention mildly spreads into surrounding connective tissue, possibly due to early-stage pseudo label noise from the teacher branch. Similar issues arise in some LYM samples, suggesting that pseudo supervision consistency remains challenging in sparse lesion scenarios.
Nevertheless, the alternating training strategy and hybrid supervision in DSAGL help mitigate attention drift. By iteratively optimizing the teacher and student branches, the model achieves more stable learning and produces class-relevant attention patterns.
5 Discussion
The proposed DSAGL framework exhibits several strengths in the context of weakly supervised WSI classification. By introducing a dual-branch architecture, the model effectively narrows the supervision gap between bag-level labels and instance-level learning. The alternating training strategy facilitates bidirectional knowledge transfer between the teacher and student branches, enhancing semantic consistency and training stability. Furthermore, the integration of a Mamba-based encoder enables efficient modeling of long-range dependencies across image patches, while maintaining linear computational complexity. Experimental results across multiple datasets confirm that DSAGL achieves superior performance compared to state-of-the-art MIL methods, particularly in terms of instance-level precision and boundary localization accuracy.
Nevertheless, several limitations remain. First, in scenarios with highly imbalanced patch distributions or extremely sparse positive regions, the pseudo labels generated by the teacher branch may be noisy during early training stages, potentially affecting the convergence of the student branch. Second, the current Mamba encoder treats patch sequences as one-dimensional inputs, which may overlook crucial two-dimensional spatial structures and anatomical priors inherent in histopathological images. Third, DSAGL does not explicitly model pseudo-label uncertainty, which could reduce robustness in regions with ambiguous or noisy annotations.
Overall, DSAGL is well suited for high-resolution, structurally complex, and weakly annotated pathological images, and demonstrates strong generalization across diverse data distributions. Future work will focus on enhancing the dual-branch learning paradigm by designing more adaptive feature extractors and uncertainty-aware training objectives to further improve model robustness and interpretability.
6 Conclusion
In this work, we propose DSAGL, a novel dual-branch teacher–student framework for weakly supervised whole-slide image (WSI) classification. By integrating attention-guided pseudo labeling and an alternating training strategy, DSAGL effectively bridges the gap between coarse-grained supervision and fine-grained instance-level learning. The incorporation of a Mamba-based encoder allows efficient modeling of long-range dependencies with linear complexity, enabling the model to process large-scale instance sets while preserving global context. Additionally, the FASA module enhances the model’s ability to focus on sparse but diagnostically significant regions through multi-scale attention. Experimental results on three benchmark datasets demonstrate that DSAGL consistently outperforms existing MIL-based approaches in both instance-level and bag-level classification tasks, particularly under challenging conditions such as low positive-instance ratios and heterogeneous lesion patterns. These findings highlight the effectiveness, scalability, and generalization potential of DSAGL, making it a promising solution for real-world computational pathology applications.
[…]
Acknowledgments
This work was supported by the 2024 College Students’ Innovative Entrepreneurial Training Plan Program, Program Number: 20240151.